papijay:

What is the value of s in the equation 3 r equals 10 plus 5 s, when r equals 10?

Gdeinward:

Well first, you must write the equation out. 3r= 10+5s

Gdeinward:

so if r equals ten, replace any r variables with 10 and simplify

Gdeinward:

3 * 10 = 10 + 5s 30 = 10 + 5s

Gdeinward:

Now we must isolate the s variable to find the answer: 30 = 10 + 5s 30 - 10 = 10 - 10 + 5s 20 = 5s 20/5 = 5s/5 4 = s